OpenStudy (anonymous):

Trouble with eigen value problems. I don't get how to do these when the given conditions are on different derivatives, i.e. y(0) and y'(pi). Find all real eigen values and eigen functions for the following problem: y'' + (r-2)y = 0 y(0)=y'(pi)=0

OpenStudy (anonymous):

r is a constant?

OpenStudy (anonymous):

\[y''=-k^2y\]\[y=(A+iB)e^{ikt}\]\[y=(Acos(kt)-Bsin(kt))\]\[y=(Acos(\sqrt{2-r}t)-Bsin(\sqrt{2-r}t))\] \[y(0)=0\] \[ \Rightarrow (Acos(\sqrt{2-r}t0)-Bsin(\sqrt{2-r}0))=A=0\]\[\Rightarrow A=0\]

OpenStudy (anonymous):

\[y'=(-A\sqrt{2-r}sin(\sqrt{2-r}t)-B\sqrt{2-r} cos(\sqrt{2-r}t))\] \[y'(0)=0\] \[ \Rightarrow (-A\sqrt{2-r}sin(\sqrt{2-r}0)-B\sqrt{2-r} cos(\sqrt{2-r}0))=-B\sqrt{2-r}=0\]\[\Rightarrow B\sqrt{2-r}=0\]

OpenStudy (anonymous):

in the problem the r was actually a lambda. i rewrote it as r for ease in typing! thanks for the continued help =]

OpenStudy (anonymous):

Is that the format of the answer you were looking for?

OpenStudy (anonymous):

in the examples that my professor has done in class, there has always been a 3 case approach to solving these problems: when r=0 ; r >0 ; and r<0. these cases will either produce trivial solutions or eigenvalues.

OpenStudy (anonymous):

it seems you only performed one case. did you just not show the others because they were trivial?

OpenStudy (anonymous):

I think all cases are contained in \[B \sqrt{2-r}=0\] Either r=2 and B=anything or B=0 and r=anything r cannot be >2 as you asked for real solutions

OpenStudy (anonymous):

I'm not sure which counts as trivial (probably B=0, as then y=0)

OpenStudy (anonymous):

is there any reason you did not use the second given condition, y'(pi)=0?

OpenStudy (anonymous):

Sorry, I misread that as y'(0)=0

OpenStudy (anonymous):

It doesn't affect it though

OpenStudy (anonymous):

as sin(pi)=sin(0)=0 and cos (pi)=-cos(0)

OpenStudy (anonymous):

that makes sense to me now. i am having a bit of trouble understanding how you got to your first step tho. would you mind if i uploaded a photo of how i usually approach these problems?

OpenStudy (anonymous):

I defined it like that to make the DE easier

OpenStudy (anonymous):

as you can see, its usually a pretty lengthy process. im just not sure how you were able to solve it so fast

OpenStudy (anonymous):

no i did not. i believe in this problem the variable r could be used to represent x if wanted. when simplified, the problem comes to (xy')' + ry - 2y = 0

OpenStudy (anonymous):

So is r not a constant?

OpenStudy (anonymous):

i dont believe so. the questions does not explicitely state it as such, so i would assume it to be a variable

OpenStudy (anonymous):

Sorry, I assumed this

OpenStudy (anonymous):

oh dear...

OpenStudy (anonymous):

no problem! do you know how to go about it with it being a variable?